OpenStudy (anonymous):

A map is drawn to a scale at 1/30000 find the actual area of a pond in km^2 if it's map area is 60cm^2

OpenStudy (anonymous):

Just use proportion. map area =1 then actual = 30000 map area = 60cm^2 thn actual = 30000*60=1800000cm^2 Just check number of 0's To convert cm^2 to km^2 1cm=1/10000 km 1cm^2=1/10^10 km^2

OpenStudy (anonymous):

I think I get it, but could you write down steps just in case? Sorry, slow learner

OpenStudy (anonymous):

I think the steps i've written would be sufficient, i can't think of anything else :)

OpenStudy (anonymous):

Alright, alright, thanks